A cannon ball barrel is 6m long. The muzzle velocity of a bullet from the cannon is 305 m/s. What is the average acceleration of the bullet going through the bullet?
1
Expert's answer
2020-01-30T09:36:23-0500

Assuming that the acceleration is constant will write the equation for displacement (1) and acceleration (2):


"s = s_0 + v_0t + \cfrac{at^2}2 \qquad \text{(1)} \\

a = \cfrac{v-v_0}t \qquad \text{(2)}"


Сonsidering that the initial velocity ("v_0" ) and initial displacement ("s_0" ) are zero, can write:


"s = \cfrac{at^2}2 \qquad \text{(1.a)} \\

a = \cfrac v t \qquad \text{(2.a)}"


Multiply (1.a) of "2a" :


"2as = a^2t^2 \qquad \text{(1.b)} \\

a = \cfrac {v} {t} \qquad \text{(2.b)}"


Substitute the expression for "a" from (2.b) into the right side of (1.b)


"2as = \Big(\cfrac v t\Big)^2 t^2 = v^2 \qquad \text{(1.c)}"


Define the acceleration:


"a = \cfrac {v^2} {2s} = \cfrac {305^2} {2*6} \approx 7752 \text{ (m/s)}"